OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = (x – 3)(x + 1)(x + 5)?

OpenStudy (abb0t):

set them all equal to zero and solve for x. (x-3) = 0 (x+1) = 0 (x+5) = 0

OpenStudy (anonymous):

3,-1,-5

OpenStudy (anonymous):

Yep =) x = 3 x = -1 x = -5 You're right! ^.^